A candy store sells malted milk balls for $2.50 per pound, and chocolate-covered raisins for $3.75 per pound. How many pounds of each should be used to make a 10 pound mixture that sells for $3.00 per pound?


Answer by stanbon(75887)

A candy store sells malted milk balls for $2.50 per pound, and chocolate-covered raisins for $3.75 per pound. How many pounds of each should be used to make a 10 pound mixture that sells for $3.00 per pound?
MM Ball DATA:
amount=x lbs ; value=2.5x dollars
CC Raisin DATA:
amount=10-x lbs ; value= 3.75(10-x)= 37.5-3.75x dollars
Mix DATA:
amount=10 lbs ; value= 3(10)= 30 dollars
EQUATION:
value + value = value
2.5x + 37.5-3.75x = 30
-3.75x=-10
x= 2.667 lbs (amount of MM Balls)
10-x=7.333 lbs (amount of CC Raisins)
Cheers,
Stan H.